OpenStudy (anonymous):

In a certain town, 20% of people commute to work by bus. If a person is selected randomly from the town, what are the odds against selecting someone who commutes by bus?

OpenStudy (shane_b):

If 20% commute by bus, the odds are 80% against selecting someone who commutes by bus. Simplifying it the odds would be\[\frac{80}{20}=\frac{4}{1}\]So you are 4 times more likely to select a person who does not commute by bus.

OpenStudy (anonymous):

in general will the will the percent that commutes be the numerator??

OpenStudy (anonymous):

and how did u het 4/1 from 80/20

OpenStudy (shane_b):

Since the question states "odds against" the chance of picking a non-bus commuter should be in the numerator. If it said "what are the odds in favor of selecting a bus commuter it would be 20/80...or 1 in 4. 80/20 = 4/1 :)

OpenStudy (shane_b):

I just reduced the fraction

OpenStudy (anonymous):

ok i see:) thanks again Shane !!!

OpenStudy (shane_b):

Wait...I goofed.

OpenStudy (anonymous):

really??lol do tell...

OpenStudy (shane_b):

Well, I sort of goofed. The final odds against would be 4 out of 5 (80%)....and odds in favor would be 1 in 5 (20%). Brainfart.

OpenStudy (anonymous):

why 4 out of 5??? pls explain

OpenStudy (shane_b):

\[80 \% =\frac{80}{100}=\frac{4}{5}\]\[20 \% =\frac{20}{100}=\frac{1}{5}\]

OpenStudy (anonymous):

oh so 80% is thwhat is left from the 20% that do take the bus??

OpenStudy (shane_b):

Yes

OpenStudy (shane_b):

The odds are still 4:1 against picking a bus-commuter...but the conclusion is that you have an 80% chance of doing that (4 out of 5).

OpenStudy (shane_b):

Sorry if I confused the heck out of you :)

OpenStudy (shane_b):

Maybe this will make it more clear. A ratio of 4:1 is the same as \[\frac{4}{total \space possibilities}=\frac{4}{4+1}\]

OpenStudy (anonymous):

so those are both options for answers 4:1 ,and 4:5?

OpenStudy (shane_b):

The odds are 4 to 1 against...which is equivalent to saying there's an 80% chance against. I'm not sure what form of the answer is expected in your problem though.

OpenStudy (anonymous):

oh ok i see thank you so much for clarifying

OpenStudy (shane_b):

Clear as mud, I know :) I need to work on being more clear from the beginning when answering these questions!

OpenStudy (anonymous):

nah your good 10 x better than any tutor at school lol

OpenStudy (shane_b):

thanks, good luck!